# NeuroPictor: Refining fMRI-to-Image Reconstruction via Multi-individual Pretraining and Multi-level Modulation

### v1

@@ Abstract
Recent fMRI-to-image approaches mainly focused on associating fMRI signals with specific conditions of pre-trained diffusion models. These approaches, while producing high-quality images, capture only a limited aspect of the complex information in fMRI signals and offer little detailed control over image creation. In contrast, this paper proposes to directly modulate the generation process of diffusion models using fMRI signals. Our approach, NeuroPictor, divides the fMRI-to-image process into three steps: i) fMRI calibrated-encoding, to tackle multi-individual pre-training for a shared latent space to minimize individual difference and enable the subsequent cross-subject training; ii) fMRI-to-image cross-subject pre-training, perceptually learning to guide diffusion model with high- and low-level conditions across different individuals; iii) fMRI-to-image single-subject refining, similar with step ii but focus on adapting to particular individual. NeuroPictor extracts high-level semantic features from fMRI signals that characterizing the visual stimulus and incrementally fine-tunes the diffusion model with a low-level manipulation network to provide precise structural instructions. By training with over 60,000 fMRI-image pairs from various individuals, our model enjoys superior fMRI-to-image decoding capacity, particularly in the within-subject setting, as evidenced in benchmark datasets. Project page: https://jingyanghuo.github.io/neuropictor/.
@@ 1 Introduction
Furthermore, in the "sorted" stage, researchers often employ latent codes extracted from fMRI signals to guide image generation through the Stable Diffusion (SD) model. Nevertheless, these codes tend to prioritize high-level semantics over low-level details, as fMRI voxels are aligned with CLIP text [29] and image features [38] to capture semantic information. However, using latent codes directly for the SD presents challenges due to the fundamental mismatch between the fMRI encoder/decoder for 1D signals and the SD for 2D images. Consequently, imperfections in both the fMRI signal-to-latent and latent-to-image projections lead to inevitable information loss and error accumulation, resulting in blurry reconstructions with lacking clear structure and pose information.
@@
Formally, our NeuroPictor initially learns a universal fMRI latent space across individuals through multi-individual pre-training to capture all original signal information and individual differences. We use an fMRI encoder-decoder architecture to reconstruct the fMRI signal as 2D brain activation flatmaps. Afterward, we utilize the encoder to map the original fMRI signal into latent codes. We introduce a High-Level Guiding Network (HLGN) to extract semantic-related signals essential in fMRI representations. The HLGN includes dual encoders aligning with text information and providing supplementary data to enhance generative model adaptability. Additionally, we incorporate a Low-Level Manipulation Network (LLMN) to collaborate with the diffusion model by directly manipulating underlying features. Recent studies like [49, 26] have shown advancements through feature map adjustments in generative models to control fine-grained visual details. Our LLMN combines a feature transformation technique with a U-net Encoder to guide image generation by updating feature maps. By decoupling the decoding process and separately controlling high-level semantics and low-level manipulation, our NeuroPictor achieves precise fMRI-to-image reconstruction.
We propose to learn fMRI-to-image decoding in three steps. i) The fMRI calibrated-encoding step. We pre-train the fMRI signal processing model for fMRI reconstruction, calibrating fMRI signals of different individuals to a shared latent space. ii) The fMRI-to-image cross-subject pre-training step. We pre-train our full model directly for fMRI-to-image decoding, leveraging over 60k fMRI-image pairs from different individuals. iii) For particular individual, we perform single-subject refining, using the same training strategy of step ii but focues on the refinement on the single subject. Our results demonstrate the benefits of multi-individual pre-training for within-subject image reconstruction and the potential of the foundation fMRI decoding model. Extensive experiments validate the efficacy of our NeuroPictor.
Contribution. Our contributions are listed as follows: (1) We propose pre-training on multiple-individual fMRI-image pairs to enhence fMRI-to-image decoding for single individuals; (2) We capture diverse neural signal information and divide it into high-level and low-level guidance to supervise the diffusion generation process; (3) We integrate all the components as a unified model for cross-subject pre-training and single-individual refining, adapting the full model directly based on the supervision of fMRI-to-image decoding.
@@ 2 Related Work
fMRI Representation. Learning a cross-subject representation for the fMRI data could enhance the down-stream tasks. In [7, 9], the masked brain modeling is pretrained on the Brain, Object, Landscape Dataset (BOLD5000) [6] to grasp useful context knowledge of fMRI data. However, due to the varied lengths of flattened 1-D fMRI voxels across individuals, only a single model can be trained per subject for subsequent image decoding. Recently, Qian et al. [31] present a large-scale transformer-based fMRI autoencoder using the UK Biobank dataset (UKB) [25], transforming individual native-space fMRI signals into unified 2D brain activation images for multi-individual brain modeling in a large-scale latent space. Different from previous works [7, 31] learn an individual-specific fMRI-to-image decoder, our NeuroPictor for the first time shows that leveraging the pretraining on multiple-individual fMRI-image pairs, can actually significantly improve the fMRI-to-image decoding of one individual.
@@ 3 Method
Figure 3: Our NeuroPictor framework is trained in three steps for fMRI-to-image decoding. i) the fMRI calibrated-encoding stage, which establishes a universal latent fMRI space across multiple individuals; ii) the fMRI-to-image cross-subject pre-training stage, which achieves multi-level modulation through perceptual learning. iii) the fMRI-to-image single-subject refining stage, using the same strategy in step ii but focuses on refinement on particular subject. The low-level manipulation network and high-level guiding network learn high-level semantics and low-level structures conditions, respectively. Some information, indicated by yellow dashed lines, is only utilized for training.
Overview. The framework of NeuroPictor is illustrated in Fig. 3. NeuroPictor is trained by three steps of fMRI calibrated-encoding, cross-subject pre-training and single-subject refining for fMRI-to-image decoding.
@@
ii) In the cross-subject pre-training stage, we learn to extract high-level semantic features (Sec. 3.2) and low-level structure features (Sec. 3.3) from the above learned fMRI latent, and then use these features as conditions to guide the generation process of diffusion model. These components are integrated via a unified model to directly train on fMRI-to-Image decoding task on multi-subject fMRI-image pairs.
iii) For the particular new single-subject, we use the same training strategy as in step ii but focus on the single subject to refine our NeuroPictor.
@@ 3.2 High-Level Semantic Feature Learning
Figure 4: Framework of our High-Level Guiding Network and Low-Level Manipulation Network
@@
However, the text feature $\bm{F}^{txt}$ extracted from the text caption naturally lacks some semantic information found in images that may not be explicitly described in the captions. Therefore, directly using the hard-mapped feature $\bm{\tilde{F}}^{txt}$ is not ideal. To address inaccurate semantic information learned in the fMRI-to-text branch, we introduce an auxiliary encoder to obtain an auxiliary semantic code $\bm{F}^{au}$. The architecture of the auxiliary encoder resembles that of the fMRI-to-text encoder, with the addition of a 1D zero convolutional layer in the final layer to initialize $\bm{F}^{au}$ as a zero tensor. This auxiliary semantic code is then residually added to the text code, resulting in the semantic representation $\bm{F}^{sem}$.
@@ 3.3 Low-Level Manipulation Network
To directly manipulate feature maps in SD’s U-net, we introduce a feature transformation technique that simultaneously performs channel-wise and dimensional-wise feature learning on fMRI embeddings. This connects fMRI representation learning with feature map learning, facilitating the subsequent manipulation process. Specifically, we employ two layers of 1D convolution to handle channel-wise correlations of fMRI embeddings and the MLP to conduct dimensional-wise feature learning. Finally, we use a reshape operation to form a feature map $\bm{F}^{l}_{0}\in\mathbb{R}^{c\times h\times w}$.
@@ 3.4 Training and Inference
$\vspace{-3pt}\mathcal{L}_{dif}=\mathbb{E}_{\bm{z}_{0},\bm{t},\bm{S}^{r},\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\epsilon_{\theta}(\bm{z}_{t},\bm{t},\bm{S}^{r},\bm{F}^{sem}))\|_{2}^{2}\Big{]},$ (8)
@@
Training Strategy. Unlike previous works [7, 31] focusing on individual-specific fMRI-to-image decoders, our NeuroPictor demonstrates the significant enhancement of fMRI-to-image decoding for single individuals through leveraging pretraining on multiple-individual fMRI-image pairs. Particularly, benefiting from the unified fMRI representation in Sec. 3.1, NeuroPictor can be trained using fMRI-image pairs from different individuals. Thus, we split the training into three stages: i) fMRI calibrated-encoding: We pre-train the calibrated-encoder to enable the cross-subject fMRI latent space; ii) Multi-subject pre-training: During pretraining, the Stable Diffusion branch and LLMN are initialized with Stable Diffusion v2.1 trained on a large-scale dataset. As illustrated in Fig. 3, we jointly train the full model using the objective function in Eq. 9. We utilize approximately 64,000 fMRI-image pairs from 8 subjects in the Natural Scenes Dataset [1] for pretraining, conducting 100k iterations to obtain a generalized fMRI-to-image generation model for these 8 subjects. iii) single-individual refining: for better modeling individually independent perception patterns, we fine-tune the network for an additional 60k iterations.
@@ 4.1 Experimental setup
Dataset. We use Natural Scenes Dataset (NSD) [1] for both training and evaluation. NSD contains visual image stimulus and corresponding fMRI recordings of 8 subjects. The original images are collected from MS-COCO dataset [24], which are consisted of complex natural images. Following [28], we use the corresponding captions of the images in COCO dataset for training. For both training and evaluation, we average three trials of fMRI signal of the same images. Previous works only use the data of subjects 1, 2, 5 or 7 to train a model for each subject. Because these subjects complete all scanning sessions, sharing the same $982$ images as testing data. As our unified fMRI representations enable training across different person, we use all training data of 8 subjects to conduct a pretraining for modeling shared activity of all subjects. This scales up our training data to about 64k pairs. Then, following the setup of previous work, we finetune our NeuroPictor using the train split of subject 1, 2, 5 and 7, respectively, and evaluate our model using the shared test split.
Evaluation Metrics. We follow the metrics of Mindeye [38] to evaluate both the high-level and low-level consistency. On the low level aspect, four metrics are introduced, including pixelwise correlation, Structural Similarity Index Metric (SSIM) [46], AlexNet(2) and AlexNet(5). Pixelwise correlation and SSIM are calculated by average the correlation distance of every paired grond truth image and reconstruction, while AlexNet(2) and AlexNet(5) represent the accuracy of two-way classification using the feature of the corresponding layers of AlexNet. High level metrics are calculated by extracting the feature using specific networks. Similarly, EffNet-B and SwAV is obtained by calculating the average distance of the features exacted by EfficientNet-B1 [44] and SwAV-ResNet50 [5]. The metric for Inception [42] and CLIP [32] represent the accuracy of two-way classification based on corresponding feature.
Implementation Details. The pretraining experiments are conducted using 4 NVIDIA RTX A6000 GPUs, with a batch size set to 64. The entire network is trained for 100,000 iterations to achieve convergence. In the finetuning stage, 2 NVIDIA RTX A6000 GPUs are employed to finetune the model for each subject, spanning 60,000 iterations. The batch size for finetuning is set to 32. For both pretraining and finetuning, the AdamW optimizer is utilized with a learning rate of $10^{-4}$ and a weight decay of 0.01. In Eq. 9, the parameter $\lambda$ is set to $0.1$. Additionally, we enables NeuroPictor to adapt to Classifier-Free Guidance [20] by randomly replacing 5% of the semantic features with unconditional CLIP embeddings corresponding to empty characters in the end of the finetuning. The Stable Dissusion v2.1 serves as the initialization for SD branch. During evaluation, the unconditional guidance scale is set to 5.0.
@@ 4.2 Main Results
Table 2: Ablation study on Subject-1. "Random init" means the fMRI Encoder starts without pre-trained weights from Sec. 3.1. Instead, it’s trained from scratch using a standard Xavier uniform initialization. "Frozen" means the fMRI Encoder remains unchanged during fMRI-to-image training. "Full" initializes the fMRI Encoder with pre-trained weights, making it trainable during fMRI-to-image training. "Multi Pret." indicates whether fMRI-to-image cross-subject pre-training (subjects 1-8) is done. If not, the entire network is trained directly on Subject-1 data. The last row shows our full proposed model.
@@
In this section, we conduct a comprehensive evaluation of our NeuroPictor against existing methods, incorporating both quantitative and qualitative comparisons.
@@
Quantitative results. The results of quantitative comparison of our NeuroPictor and previous state-of-the-art methods are displayed in Tab. 1. We generate one sample for each subject and calculate the low-level and high-level metrics following the official code of [38]. We average the metrics across 4 subjects (i.e. Subject-1, 2, 5 and 7) and report both the pre-trained (i.e. “Pretrain” in line 7 of Tab. 1) and fine-tuned results. The results show that our NeuroPictor outperforms other methods in six out of eight metrics, demonstrating that our reconstruction results can maintain semantic consistency while preserving low-level details present in the original images. We find that our method falls behind in pixel correlation metrics compared to three other models. This could potentially be attributed to the fact that other methods employ an image-to-image diffusion model and use a blurred image produced by VAE as an initialization. As reported in [33, 13], such a blurred image inherently exhibits higher pixel correlation than finely reconstructed results. Furthermore, our non-finetuned model achieves results comparable to other models, except for Mindeye. This validates that our training on multi-individual data effectively captures shared perceptual features across different individuals.

### v2

@@ Abstract
Recent fMRI-to-image approaches mainly focused on associating fMRI signals with specific conditions of pre-trained diffusion models. These approaches, while producing high-quality images, capture only a limited aspect of the complex information in fMRI signals and offer little detailed control over image creation. In contrast, this paper proposes to directly modulate the generation process of diffusion models using fMRI signals. Our approach, NeuroPictor, divides the fMRI-to-image process into three steps: i) fMRI calibrated-encoding, to tackle multi-individual pre-training for a shared latent space to minimize individual difference and enable the subsequent multi-subject training; ii) fMRI-to-image multi-subject pre-training, perceptually learning to guide diffusion model with high- and low-level conditions across different individuals; iii) fMRI-to-image single-subject refining, similar with step ii but focus on adapting to particular individual. NeuroPictor extracts high-level semantic features from fMRI signals that characterizing the visual stimulus and incrementally fine-tunes the diffusion model with a low-level manipulation network to provide precise structural instructions. By training with about 67,000 fMRI-image pairs from various individuals, our model enjoys superior fMRI-to-image decoding capacity, particularly in the within-subject setting, as evidenced in benchmark datasets. Our code and model are available at https://jingyanghuo.github.io/neuropictor/.
@@ 1 Introduction
Furthermore, in the "sorted" stage, researchers often employ latent codes extracted from fMRI signals to guide image generation through the Stable Diffusion (SD) model. Nevertheless, these codes tend to prioritize high-level semantics over low-level details, as fMRI voxels are aligned with CLIP text [29] and image features [38] to capture semantic information. However, using latent codes directly for SD presents challenges due to the fundamental mismatch between the separately trained fMRI encoder/decoder for 1D signals and the SD model for 2D images. Consequently, imperfections in both the fMRI signal-to-latent and latent-to-image projections lead to inevitable information loss and error accumulation, resulting in blurry reconstructions with lacking clear structure and pose information.
@@
Formally, our NeuroPictor initially learns a universal fMRI latent space across individuals through multi-individual pre-training to capture all original signal information and individual differences. We use an fMRI encoder-decoder architecture to reconstruct the fMRI signal as 2D brain activation flatmaps. Afterward, we utilize the encoder to map the fMRI signal into latent codes. We introduce a High-Level Guiding Network (HLGN) to extract semantic-related signals essential in fMRI representations. The HLGN includes dual encoders aligning with text information and providing supplementary data to enhance generative model adaptability. Additionally, we incorporate a Low-Level Manipulation Network (LLMN) to collaborate with the diffusion model by directly manipulating underlying features. Recent studies [50, 26] have shown advancements through feature map adjustments in generative models to control fine-grained visual details. Our LLMN combines a feature transformation technique with a U-net Encoder to guide image generation by updating feature maps. By decoupling the decoding process and separately controlling high-level semantics and low-level manipulation, our NeuroPictor achieves precise fMRI-to-image reconstruction.
We propose to learn fMRI-to-image decoding in three steps. i) The fMRI calibrated-encoding step. We pre-train the fMRI signal processing model for fMRI reconstruction, calibrating fMRI signals of different individuals to a shared latent space. ii) The fMRI-to-image multi-subject pre-training step. We pre-train our full model directly for fMRI-to-image decoding, leveraging about 67k fMRI-image pairs from different individuals. iii) For particular individual, we perform single-subject refining, using the same training strategy of step ii but focues on the refinement on the single subject. Our results demonstrate the benefits of multi-individual pre-training for within-subject image reconstruction and the potential of the foundation fMRI decoding model. Extensive experiments validate the efficacy of our NeuroPictor.
Contribution. Our contributions are listed as follows: (1) We propose pre-training on multiple-individual fMRI-image pairs to enhence fMRI-to-image decoding for single individuals; (2) We capture diverse neural signal information and divide it into high-level and low-level guidance to supervise the diffusion generation process; (3) We integrate all the components as a unified model for multi-subject pre-training and single-individual refining, adapting the full model directly based on the supervision of fMRI-to-image decoding.
@@ 2 Related Work
fMRI Representation. Learning a cross-subject representation for the fMRI data could enhance the down-stream tasks. In [7, 9], the masked brain modeling is pretrained on the Brain, Object, Landscape Dataset (BOLD5000) [6] to grasp useful context knowledge of fMRI data. However, due to the varied lengths of flattened 1-D fMRI voxels across individuals, only a single model can be trained per subject for subsequent image decoding. Recently, Qian et al. [31] present a large-scale transformer-based fMRI autoencoder using the UK Biobank dataset (UKB) [25], transforming individual native-space fMRI signals into unified 2D brain activation images for multi-individual brain modeling in a large-scale latent space. Different from previous works [7, 31] learn an individual-specific fMRI-to-image decoder, our NeuroPictor shows that leveraging the pretraining on multiple-individual fMRI-image pairs, can actually significantly improve the fMRI-to-image decoding of one individual. Concurrent work [39] also aims to develop a shared-subject fMRI-to-image model. Different from their approach, which uses subject-specific ridge regression to map individuals’ fMRI data into a common space, our method directly transforms native-space fMRI signals from any individual into unified 2D activation images. This eliminates the need for subject-specific encoding modules.
@@ 3 Method
Figure 3: Our NeuroPictor framework is trained in three steps for fMRI-to-image decoding. i) the fMRI calibrated-encoding stage, which establishes a universal latent fMRI space across multiple individuals; ii) the fMRI-to-image multi-subject pre-training stage, which achieves multi-level modulation through perceptual learning. iii) the fMRI-to-image single-subject refining stage, using the same strategy in step ii but focuses on refinement on particular subject.
Overview. The framework of NeuroPictor is illustrated in Fig. 3. NeuroPictor is trained by three steps of fMRI calibrated-encoding, multi-subject pre-training and single-subject refining for fMRI-to-image decoding.
@@
ii) In the multi-subject pre-training stage, we learn to extract high-level semantic features (Sec. 3.2) and low-level structure features (Sec. 1.5) from the above learned fMRI latent, and then use these features as conditions to guide the generation process of diffusion model. These components are integrated via a unified model to directly train on fMRI-to-Image decoding task on multi-subject fMRI-image pairs.
iii) For the particular single-subject, we use the same training strategy as in step ii but focus on the single subject to refine our NeuroPictor.
@@ 3.2 High-Level Semantic Feature Learning
Figure 4: Framework of our High-Level Guiding Network and Low-Level Manipulation Network.
@@
However, directly using the hard-mapped feature $\bm{\tilde{F}}^{txt}$ is not optimal. To address inaccurate semantic information learned in the fMRI-to-text branch, we introduce an auxiliary encoder to obtain an auxiliary semantic code $\bm{F}^{au}$. The architecture of the auxiliary encoder resembles that of the fMRI-to-text encoder, with the addition of a 1D zero convolutional layer in the final layer to initialize $\bm{F}^{au}$ as a zero tensor. This auxiliary semantic code is then residually added to the text code, resulting in the semantic representation $\bm{F}^{sem}$.
@@ 3.3 Low-Level Manipulation Network
To directly manipulate feature maps in SD’s U-net, we introduce a feature transformation technique that simultaneously performs channel-wise and dimensional-wise feature learning on fMRI embeddings 222 We use the terms ”channel” and ”dimension” to refer to the second and third dimensions of the fMRI latent tensor. . This connects fMRI representation learning with feature map learning, facilitating the subsequent manipulation process. Specifically, we employ two layers of 1D convolution to handle channel-wise correlations of fMRI embeddings and the MLP to conduct dimensional-wise feature learning. Finally, we use a reshape operation to form a feature map $\bm{F}^{l}_{0}\in\mathbb{R}^{c\times h\times w}$.
@@ 3.4 Training and Inference
$\mathcal{L}_{dif}=\mathbb{E}_{\bm{z}_{0},\bm{t},\bm{S}^{r},\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\epsilon_{\theta}(\bm{z}_{t},\bm{t},\bm{S}^{r},\bm{F}^{sem}))\|_{2}^{2}\Big{]},$ (8)
@@
Training Strategy. Benefiting from the unified fMRI representation in Sec. 3.1, NeuroPictor can be trained using fMRI-image pairs from different individuals. Thus, we split the training into three stages: i) fMRI calibrated-encoding: We pre-train the calibrated-encoder to enable the cross-subject fMRI latent space; ii) Multi-subject pre-training: During pretraining, the Stable Diffusion branch and LLMN are initialized with Stable Diffusion v2.1 trained on a large-scale dataset. We jointly train the full model, excluding the auxiliary encoder, using about 67,000 fMRI-image pairs from 8 subjects in the Natural Scenes Dataset [1] for pretraining, conducting 100k iterations to obtain a generalized fMRI-to-image generation model for these 8 subjects. iii) single-individual refining: for better modeling individually independent perception patterns, we fine-tune the network for an additional 60k iterations.
@@ 4.1 Experimental setup
Dataset. We use Natural Scenes Dataset (NSD) [1] for both training and evaluation. NSD contains visual image stimulus and corresponding fMRI recordings of 8 subjects, with each subject viewing 8,000-9,000 images. The original images are collected from MS-COCO dataset [24], which are consisted of complex natural images. Following [28], we use the corresponding captions of the images in COCO dataset for training. For both training and evaluation, we average three trials of fMRI signal of the same images. Previous works only use the data of subjects 1, 2, 5 or 7 to train a model for each subject. Because these subjects complete all scanning sessions, sharing the same $982$ images as testing data. As our unified fMRI representations enable training across different person, we use all training data of 8 subjects to conduct a pretraining for modeling shared activity of all subjects. This scales up our training data to about 67k pairs. Then, following the setup of previous work, we finetune our NeuroPictor using the train split of subject 1, 2, 5 and 7, respectively, and evaluate our model using the shared test split.
Evaluation Metrics. We follow the metrics of Mindeye [38] to evaluate both high-level and low-level consistency. On the low-level aspect, we use pixelwise correlation, Structural Similarity Index Metric (SSIM) [47], AlexNet(2), and AlexNet(5). High-level metrics are calculated by extracting features using specific networks, including EffNet-B [45], SwAV [5], Inception [43], and CLIP [32]. Please refer to Supplementary for more details.
Implementation Details. The pretraining experiments are conducted using 6 NVIDIA RTX A6000 GPUs, with a batch size set to 96. During this stage, we exclude the auxiliary encoder and semantic loss to learn a base model. The entire network is trained for 100,000 iterations to achieve convergence. In the finetuning stage, 2 NVIDIA RTX A6000 GPUs are employed to finetune the model for each subject, spanning 60,000 iterations. The batch size for finetuning is set to 32. In Eq. 9, the parameter $\lambda$ is set to $0.1$. Additionally, we enables NeuroPictor to adapt to Classifier-Free Guidance [20] by randomly replacing 5% of the semantic features with unconditional CLIP embeddings corresponding to empty characters in the end of each stage. During evaluation, the unconditional guidance scale is set to 5.0.
@@ 4.2 Main Results
Table 2: Ablation study on Subject-1. "Random init" means the fMRI Encoder starts without pre-trained weights from Sec. 3.1. Instead, it’s trained from scratch using a standard Xavier uniform initialization. "Frozen" means the fMRI Encoder remains unchanged during fMRI-to-image training. "Full" initializes the fMRI Encoder with pre-trained weights, making it trainable during fMRI-to-image training. "Multi Pret." indicates whether fMRI-to-image multi-subject pre-training (subjects 1-8) is done. If not, the entire network is trained directly on Subject-1 data.
@@
In this section, we comprehensively evaluate NeuroPictor against existing methods, using both quantitative and qualitative comparisons.
@@
Quantitative results. The results of quantitative comparison of our NeuroPictor and previous state-of-the-art methods are displayed in Tab. 1. We generate one sample for each subject and calculate the low-level and high-level metrics following the official code of [38]. We average the metrics across 4 subjects (i.e. Subject-1, 2, 5 and 7) and report both the pre-trained (i.e. “w/o ft” in line 7 of Tab. 1) and fine-tuned results. The results show that our NeuroPictor outperforms other methods in six out of eight metrics, demonstrating that our reconstruction results can maintain semantic consistency while preserving low-level details present in the original images. We find that our method falls behind in pixel correlation metrics compared to three other models. This could potentially be attributed to the fact that other methods employ an image-to-image diffusion model and use a blurred image produced by VAE as an initialization. As reported in [33, 13], such a blurred image inherently exhibits higher pixel correlation than finely reconstructed results. Furthermore, our non-finetuned model achieves results comparable to other models, except for Mindeye. This validates that our training on multi-individual data effectively captures shared perceptual features across different individuals.